OpenStudy (anonymous):

A tube of tennis balls has the following measures: 4-inch diameter and a height of 8 inches. What is the surface area of the tube of tennis balls? (Use 3.14 for π.) 213.52 in.2 301.44 in.2 226.08 in.2 125.6 in.2

OpenStudy (anonymous):

Cylinder Surface Area = 2*pi*r^2 + 2*pi*r*h Just plug and chug!

OpenStudy (ybarrap):

 Using the sketch above: The area of the time is \(\pi r^2\), the area of the sides is \(2\pi rh\) $$ \large{ SA=2\times\pi r^2+ 2\times \pi rh } $$ That's it!

OpenStudy (anonymous):

You have a diameter of 4, so half of that is the radius = 2.

OpenStudy (ybarrap):